Displaying reports 201-220 of 1079.Go to page Start 7 8 9 10 11 12 13 14 15 End
Reports until 16:53, Thursday 30 January 2025
GQuEST Output Filter Cavities
Torrey Cullen - posted 16:53, Thursday 30 January 2025 (12105)
Noise in Error Spectrum with Increasing Controller Gain

[Jeff, Torrey]

The background for this is we have noticed 10's of kHz ringing up in the timeseries of the error signal when the gain on the fast controller gets too high. On the computer app where we can control multiple instruments at once we have increased the gain and taken spectrums while the cavity is locked. The plan is to do this for the thorlabs laser and the teraxion laser and compare the two. We are waiting on an optical isolator to conduct the teraxion half of this experiment, in the mean time here is the thorlabs laser data.

The following data is for locking OFC1, with the thorlabs laser, while controlling from the computer app. thorlabslaservaryingcontrollergain.png are the results. We can see a clear increase in the noise around 20 kHz as the controller gain is increased. It will be interesting to see if a similar feature is there when using the teraxion laser. Note the controler UGFs are listed between 2-18 hz, but there are several gain increases between the controller and the laser, so they should not be taken as the UGF of the whole loop.

We still plan on taking TF's instead of this analysis. For now, the smallest excitation strength on the moku (1mV) is too large. We are working on this.

Images attached to this report
Vacuum Systems General (GQuEST)
Daniel Grass - posted 19:04, Wednesday 29 January 2025 (12104)
5/16-24 holes to attach the Top Flange of the LFC Input Vacuum Cube Tapped

[Ian, Daniel]

We tapped all 24 holes to secure the 10" flange on top of the Laser Filter Cavity (LFC) Input Vacuum Cube with the cleaned 5/16"-24 taps using some isopropanol as a lubricant. I used a 7 mm 12-point wrench. While its ~3" length made me apply more force to tap, I felt like it reduced the change of breaking a tap. We kept a 10" flange on top so that the vacuum system wasn't exposed. The taps fit within the holes of the flange. Before tapping, screws could only get 3 turns in by hand before encountering a lot of resistance. We used the tappered tap, plug tap, and bottom tap, going all the way to the bottom. The tappered tap went up 14 revolutions and the bottom tap went up 17 revolutions after tapping. The 1.75" long screws easily went into the holes after the plug tap and bottom tap. We used the bottom tap to ensure the screws wouldn't encounter any resistance when the flange is tightened. Some metal chips were generated. Some came out with the taps, some are surely in the bottom of the holes.

The TiCN coating might have been slightly removed by the end of tapping. The color might be different and the tapping may have gotten harder, but this is somewhat subjective.

Electronics General
Jeffrey Wack - posted 16:54, Wednesday 29 January 2025 (12103)
cymac main chassis - modification to ATX connector attached to power supply to power clock PCB

Context: I am following the instructions of LIGO DCC document E2300448-v2 to assemble the main chassis of our 'standard' cymac, named Babbage.

See D2300234-v1 for instuctions.

Notes:

I borrowed a tool from Todd Etzel in Downs which allows you to remove the metal connectors from the plastic ATX sheath. Without this tool it was near impossible, and I believe I made the job much harder for myself by trying to simply yank the wires out. I reccomend you do not attempt to remove wires from an ATX connector without one of these tools.

After this the process was straightforward, and described resonably well in the document above.

Images attached to this report
Vacuum Systems General
Daniel Grass - posted 12:31, Wednesday 29 January 2025 (12101)
5/16-24 TiCN coated High Speed Steel Taps Cleaned for use around UHV

I cleaned some 5/16"-24 TiCN coated High Speed Steel Taps for use around UHV. These taps are stronger and more durable than the taps in the tap set. I did a 90 second bath and sonication in Simple Green:DI water (1:30) and noticed a few particulates come off (I don't think it's the coating as the coating looked intact, but I didn't want to risk stripping it, hence only 90 seconds). I brushed the taps, washed with DI water, did a 2 minute sonication in a DI water bath, washed with isopropanol, then did a 2 minute sonication in isopropanol.

GQuEST Output Filter Cavities
Torrey Cullen - posted 16:38, Tuesday 28 January 2025 - last comment - 15:10, Wednesday 29 January 2025(12099)
Lab Work 1/28 - Continued OFC3 Work

[Jeff,Torrey]

Assorted notes:

-I completed the assembly of the piezo mirror. This required having daniel machine a custom spacer for the assembly. With the spacer, the piezo alignment tool works. Note that this new piezo has a larger OD that is large enough to prohibit the use of the nylon tipped set screws for alignment. Because of this the spacer and piezo may not be concentrically aligned to the mirror.

-Installed a 2 meter 780 PM fiber at the OFC2/OFC3 sled fiber port hub. This allows a connection from the third AOM path to OFC3 775 input.

-We are going to start mode matching into cavity 3. The USB extension from the NUC computers does not recognize the scanning slit beam profiler. This is probably due to the USB extension being over 100 ft and needing external power. We are going to use the new lab laptop for now instead.

-The new laptop is unusable for the scanning slit software. For some reason it is constantly crashing and freezing.

-After switching to Torrey's laptop, we collected data for the beam size between the launcher and the third filter cavity. We noticed amplitude fluctuations on the beam, but they reduced when we raised the power on this path (by turning the amplifier current to 1A. This should correspond to ~1.5W total output on the distribution center).

-We spent some time profiling and finding a mode matching solution. The waist for this light needs to be .26 meters after the collimator for this configuration which is making it difficult to mode match. We are going to try moving the collimator back tomorrow and reprofiling.

-We turned the amplifier back down until tomorrow.

Comments related to this report
Torrey Cullen - 12:05, Wednesday 29 January 2025 (12100)

After playing around in JAMMT I found that moving the collimator back would result in easy mode matching solutions. I moved the collimator back 12 inches and reprofiled. Now, according to finesse and JAMMT, a robust solution is f=2000 mm @ z=0.066, f=2500 mm @ z=0.354, (v=100 %). We have these lens. Will implement after lunch.

Torrey Cullen - 15:10, Wednesday 29 January 2025 (12102)

I put the two lens in the positions according to the above comment. I then profiled the beam after the second lens. This gives the following data below:

                                         Horizontal    Vertical
Beam waist radius[um]    402.531593    423.098793
Waist position [m]            0.148341    0.146998
Rayleigh range [inches]   25.859159    28.569197


Radius Average               412.815193
Waist Position Average    0.147669

According to finesse, and accounting for the extra two 1/8 spacers, the beam this cavity accepts is 411 um @ .106 meters after the lens. The measured quanity is 412 um .147 meters after the lens. This should yield a mode mismatch well below 1%. I will close this work for now unless we see large laguerre gauss modes in the cavity scans in the future.

Images attached to this comment
GQuEST Output Filter Cavities (GQuEST)
Daniel Grass - posted 14:04, Tuesday 28 January 2025 (12098)
Spacer to hold the 2 mm thick NAC 2125 Piezo Made

[Torrey, Daniel]

I designed and made a spacer to go between the top of the piezo assembly and the NAC 2125 piezo (2 mm thick) so that the mirror is within the piezo base and the piezo can be centered with the centering tool I previously made and the #4-40 nylon tipped set screws. The spacer is 0.75" OD, 0.469" ID, and 0.226" thick. I designed it to be 0.21" thick (see photo), but this came out a bit longer because I thought I would have to machine the surface to get it smooth. Instead, parting at 400 rpm gave a nice surface finish.

I measured the inside of the piezo to be ~0.468". It fit over the 0.46" diameter part of the centering tool, but not the 0.47" diameter part of the centering tool. 

Images attached to this report
GQuEST General (GQuEST)
Daniel Grass - posted 13:54, Tuesday 28 January 2025 - last comment - 10:27, Sunday 02 February 2025(12097)
Mounting Plate for the TeraXion LXM-U Laser Made

I designed a plate to mount the TeraXion LXM-U Laser to the optics table with good thermal contact. I used a fly cutter at 1800 rpm to get a really nice top and bottom surface finish and a 3/4" end mill to get the sides. See attached files.

I unfortunately broke a tap in one of the mounting holes and couldn't remove it. Martin Mendez in the Chemistry Machine Shop said it would take 2-3 hours of shop time to burn through the tap with the plunge EDM, costing $120-$190. Since only 2-3 holes are needed to securely mount the laser, I cut out the hole with the broken tap with the band saw. 

For the other 3 #4-40 holes, I drilled them out to 0.0935" (using a #42 drill) instead of the standard 0.089" (a #43 drill) to help with the tapping. This is still plenty of thread engagement, and the tapping went much better.

Non-image files attached to this report
Comments related to this report
Daniel Grass - 10:27, Sunday 02 February 2025 (12109)

I cleaned the plate with Simple Green:DI water, DI water, and isopropanol and sonication for each step. The LXM-U fits on the #4 holes, seemingly with a great contact. One of the 1/4-20 counterbored holes isn't useful as a laser input/output blocks it, but there are still 3 holes to secure the plate.

Equipment Loans
Torrey Cullen - posted 16:06, Monday 27 January 2025 - last comment - 16:19, Monday 27 January 2025(12095)
IR Viewer Return to Cryo

Update to [12045]:

The 1550 IR viewer and the additional lens attachment have been returned to the cryo lab lista cabnet in the drawer labeled "IR Viewer".

Comments related to this report
Ian MacMillan - 16:19, Monday 27 January 2025 (12096)

This post has also been posted in an Adhikari Cryo Lab elog post

Vacuum Systems General
Daniel Grass - posted 22:16, Friday 24 January 2025 (12093)
5/16-24 Thread Chaser and 7/16 in socket cleaned; LFC Input Vacuum Cube Top Flange Threads still problematic

I cleaned the 5/16"-24 thread cleaner that I previously machined and a 7/16" socket in acetone. I rinsed the paint off the 7/16" socket with the acetone spray bottle before putting it alone in a 100 mL beaker with acetone and sonicating for 3 minutes, replacing the solvent, and sonicating for another 7 minutes. I sonicated the thread chaser and socket in the same beaker with isopropanol for 10 minutes.

I used the thread chaser on a few holes on the LFC Input Vacuum Cube Top Flange, but the threads did not seem greatly improved. Either we just power through and hope the threads are actually fine with the silver coated screws, or we re-tap the threads. We have a 5/16"-24 tap, but I think we want some slightly nicer ones, including a bottom tap (which I don't think we have) to reach the bottom threads.

GQuEST Output Filter Cavities
Jeffrey Wack - posted 18:03, Friday 24 January 2025 - last comment - 15:57, Monday 27 January 2025(12092)
First contact of optics for output filter cavity number three (OFC 3)

[Torrey, Jeff]

We performed a first contact cleaning procedure on the four optics planned for use in output filter cavity three. We struggled with the first contact solution creating strings, and two of the four optics had to be re-painted and will be peeled at a future date.

For each optic, the procedure was as follows:

1) Remove the optic from the package, and place it in a mount, HR side down. Create a label for this optic.

2) Paint the AR side of the optic with a first coat. Add a strip of plastic mesh to serve as a handle on removal, and paint another coat on top.

3) Wait for the first contact on the AR side to dry completely, then place it face down on a clean surface. We used a TEX wipe 1109. Based on the fact that later the first contact was stuck to the wipe, the 5 minutes or so that we waited was not long enough.

4) Repeat step 2 on the HR side.

5) Wait for the HR side to dry completely

6) Peel off the first contact, using the tweezers to lift the mesh handle to start. Be mindful of strings! They are likely to lurk around the edges of the optic as you pull away, and they will make you very sad if they land on the optic. We used the tweezers to remove strings if they were close to the edge of the optic, but when the optic had strings on the middle we re-painted that face.

7) Place clean optics into their mounts and label the mount.

 

Images attached to this report
Comments related to this report
Torrey Cullen - 15:57, Monday 27 January 2025 (12094)

I peeled the first contact that was left on over the weekend. The mirrors seem sufficiently cleaned. The first three mirrors are in their mounts and on the table. Note that we are trying the syskew mounts instead of the newport flexture mounts this time. I am having trouble assembling the last mirror that is compressed against the piezo . We need the nylon tipped screws that hold the piezo in place while assembling. This piezo (nac2125 from CTS) has been soldered to a BNC connection for later use. I think we also had bought a custom spacer as this piezo is not as tall as the noliacs that were in the original design. I can't quite finish assembling this until we get some stuff. In the mean time, it has a SM1 ring and a viton spacer ring in place.

All of these new superoptics have been labelled and their position recorded on the wiki.

 

Images attached to this comment
Lab Infrastructure General
Daniel Grass - posted 15:09, Friday 24 January 2025 (12091)
Drill Press Chuck Re-seeded and its Platform Centered and Leveled

I put the chuck into the drill press, and it seemed to be running fairly smoothly. I also rotated the platform to make it level and rotated the platform about the vertical support column so that the platform's hole was directly below the chuck.

Vacuum Systems Pumps (GQuEST)
Daniel Grass - posted 15:07, Friday 24 January 2025 (12090)
Agilent Vacuum Pump Mostly Assembled

[Alex, Daniel]

We began to assemble the Agilent Vacuum Pump for the SNSPD Dewer. We mostly followed the manual, but needed to find longer M3 screws (~10 mm long instead of 6 mm) to mount the bracket to the turbo pump (with the fan and the lock washers). We may have damaged the threading by insterting the included screws which were not long enough, so I re-tapped them.

Seperately, next time, it will be easier to assemble the scroll pump supports first with nothing else on the base plate.

After the scroll, turbo, and controller were assembled to the base plate, we wheeled the system into the clean room. We connected the included KF16 elbow to the Agilent turbo pump and KF16 to KF25 tubing from the turbo pump to the scroll pump. They included an o-ring with a mesh filter, so we attached it between the tubing and the scroll pump.

We started to attach electronics to the pumps and controller, but there isn't room on the controller for the cables. This seems like a design oversight on Agilent's end. We can mount the controller somewhere else near the pumps/dewer.

We need to design, purchase, and assemble the tubing and other vacuum components between the turbo pump and the main vacuum chamber. This is ongoing and Alex will add a cart to Techmart/OpenProject later today.

Vacuum Systems General
Daniel Grass - posted 20:24, Thursday 23 January 2025 (12088)
5/16-24 Thread Chaser Cleaned, Tested, and Machined

I cleaned a 5/16"-24 Thread Chaser to clean the threads for CF Flanges, specifically the threads for the top flange on the Laser Filter Cavity (LFC) Input Vacuum Cube. Because the thread chaser is made of carbon steel, I cleaned it in a 10 minute acetone bath/sonication and a 10 minute isopropanol bath/sonication. I first had to reclean the acetone container with DI water/simple Green and DI water because there was residual grease/gunk in the container from cleaning the locking pliers.

I used the thread chaser on one of the screw holes with the flanges still attached, and it seemed to allow the screw to go in deeper (I didn't do a rigorous measurment, partially because it's a bit ad hoc when to stop tightening the screw). The diameter of the threader chaser by the hex head is too large to fit through a 10" CF Flange (0.96" thick) and thread 0.75" deep into the vacuum cube. To get around this issue, I used the lathe to machine down the tip of the thread chaser to hypothetically fit into a 1.58" long, 0.312" diamter (< 5/16") cylinder so that it can be inserted all the way down the threaded holes. I can use the counterbored holes (0.25" deep) on the 10" flange from the holometer to get the remaining depth needed (1.83" measured). There is still ~0.3" left on top of the thread chaser to use so that a socket can be used. The narrowest part of the thread chaser is ~1/4" thick and ~1" long, so I don't think I meaningfully weaken it (famous last words). Nevertheless, one should be gentile with it.

Tomorrow, I plan on cleaning the 7/16" socket and re-cleaning the thread chaser to see if I can fully clean up the threads.

GQuEST Output Filter Cavities
Torrey Cullen - posted 16:55, Thursday 23 January 2025 - last comment - 11:40, Friday 24 January 2025(12087)
Planning for OFC3

I've updated the layout diagram for OFCs 1, 2, and 3 for its current configuration in B111B. I've laid out all the optical elements on the table for a proof of concept. It will be cramped but it should fit. This configuration should allow continuous building and testing of OFC3 while allowing for chained cavities in the future. Couple of other notes:

-For this to be fully functional we need to buy a 775 50:50 BS and a thorlabs PDA10A2.

-Note the oxidation (I think?) in PXL_20250124_000346938MP.jpg. Unsure if we care, just pointing it out.

-Still need to calculate mode matching solutions for both 1550 inputs and the 775 inputs.

-The .svg that created this image is at "\Nextcloud\GQuEST\Layout_Mockups\readout_FCs_1_2_and_3.svg"

-It is unclear how the output of OFC3 will get to the eventual OFC4. 

Images attached to this report
Non-image files attached to this report
Comments related to this report
Torrey Cullen - 11:40, Friday 24 January 2025 (12089)

My current plan is to add two additional 1/8 inch spacers to OFC3. This should change the FSR by c/2.4 - (c/(2.4+2/8*.0254)) ~ 300kHz. With a linewidth in the 10's of kHz this should provide enough separation between cavities. This means the waist of the 1550 beam should be 691 um which can be achieved by putting a f=1m lens just after the output of OFC2 (99.92% mode overlap).

Images attached to this comment
Vacuum Systems General
Daniel Grass - posted 16:39, Thursday 23 January 2025 (12086)
Front Face of Laser Filter Cavity Input Cube Mostly Assembled

[Jeff, Daniel]

We attached a 10" to 6" reducer flange to the front face (where the laser enters) to the Laser Filter Cavity (LFC) Input cubed and tightened the screws to 34 Nm. We attached a 6" to 2x2.75" reducer flange to this reducer flange, and I tightened the screws to 34 Nm using the new 2" torque adaper extension tool. This new tool doesn't affect the torque amount if it's perpendicular to the handle.

I then attached a 2.75" blank flange to the unused port on the 6" to 2x2.75" reducer flange and tightened the screws to 14 Nm.

Displaying reports 201-220 of 1079.Go to page Start 7 8 9 10 11 12 13 14 15 End